A source emits light of wavelengths 555 nm and 600 nm. The radiant flux of the 555 nm part is 40 W and of the 600 nm part is 30 W. The relative luminosity at 600 nm is 0.6. Find (a) the total radiant flux, (b) the total luminous flux, (c) the luminous efficiency.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. The radiant flux of 555nm part is 40W and of the 600nm part is 30W (a) Total radiant flux = 40W + 30W = 70W (b) Luminous flux = (L.Fllux) base 555nm + (L.Flux) base 600nm = 1 × 40× 685 + 0.6 × 30 × 685 = 39730 lumen (c) Luminous efficiency = Total lu min ous flux/Total radiant flux = 39730/70 = 567.6 lumen/W